OpenStudy (angelicahicks29):

(1-cos^2x)(1+cos^2x)=2sin^2x-sin^4x

OpenStudy (anonymous):

1-(cos^2x)=sin^2x 1+cost^2x=1+(1-sin^2x) (2-sin^2x)(sin^2x)=(2sin^2x)-(sin^4x) 2sin^x-sin^4x=2sin^2x-sin^4x

OpenStudy (anonymous):

I'm guessing the point of that was to prove the equation?

OpenStudy (angelicahicks29):

Yes and I'm very confused!

OpenStudy (anonymous):

Ok let me explain

OpenStudy (anonymous):

sin^2x+cos^2x=1 Did you learn this equation?

OpenStudy (angelicahicks29):

Yes I figured it out with your response, I started out writing the problem that way and I felt like I was doing something wrong.